I am having trouble with this one.
2 solutions of h2so4. one is 40% concentration and one is 25% concentration. how much of each do I need to get 137 liters of 36% solution?

Found 2 solutions by stanbon, johndeere4:
Answer by stanbon(75887)

2 solutions of h2so4. one is 40% concentration and one is 25% concentration. how much of each do I need to get 137 liters of 36% solution?
------------------------
acid + acid = acid
0.40x + 0.25(137-x) = 0.36*137
Multiply thru by 100 to get:
40x + 25*137 - 25x = 36*137
15x = 11*137
x = 100.47 liters (amt. of 40% solution needed in the mixture)
137-100.47 = 36.53 liters (amt of 25% solution in the mixture)
============================================================
Cheers,
Stan H.

Answer by johndeere4(1)

Thank you so much, stanbon!!! #2418059
And thank you for #248087 as well!
And thanks to the other tutor for helping too!